Grade 11Mechanics

A train is uniformly accelerated for a distance equal to 1/m of the distance between two stations and uniformly retarded for 1/n of the distance b/w thesthese stations .find the ratio of maximum to the average velocity during the journey.
Thank for the previous answer!

Simran
8 Years agoGrade 11

1 Answer

Eshan
8 Years ago
Dear student,

Let the maximum velocity attained bev_0and acceleration for distance\dfrac{s}{m}bea_1, and that for distance\dfrac{s}{n}bea_2.
Hence we have\dfrac{s}{ m}=\dfrac{v_0^2}{2a_1}and\dfrac{s}{ n}=\dfrac{v_0^2}{2a_2}

Also the total time taken in the motion is sum of those periods with acceleration, constant velocity and retardation=t_1+t_2+t_3=\dfrac{v_0}{a_1}+\dfrac{s(1-\dfrac{1}{m}-\dfrac{1}{n})}{v_0}+\dfrac{v_0}{a_2}

Thus average velocity=\dfrac{s}{t}

Hence the required ratio=\dfrac{v_0}{s/t}=\dfrac{v_0}{s}t=\dfrac{v_0^2}{a_1s}+(1-\dfrac{1}{m}-\dfrac{1}{n})+\dfrac{v_0^2}{a_2s}

=\dfrac{2}{m}+(1-\dfrac{1}{m}-\dfrac{1}{n})+\dfrac{2}{n}=1+\dfrac{1}{m}+\dfrac{1}{n}